OpenStudy (anonymous):

Evaluate the logarithm log[3]1/9

OpenStudy (anonymous):

your job is to answer \[3^x=\frac{1}{9}\]

OpenStudy (anonymous):

if you can answer this, you can solve the problem. otherwise you are stuck more or less

OpenStudy (anonymous):

so is this possible?

OpenStudy (anonymous):

x=-2 ;)

OpenStudy (anonymous):

hipster will give you a hint. he (she) is fond of them...

OpenStudy (anonymous):

hahahahahaa this is true. the answer lies on the interval (-1,1)

OpenStudy (anonymous):

cough, 0. cough

OpenStudy (anonymous):

actually if you have a calculator and wish to cheat (aka not think) i can show you how to get the answer. it is what i would do

OpenStudy (anonymous):

do show us, satellite.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

got a calculator handy? we will find the answer (-2) quickly but first lets be clear. the answer is -2 because \[3^{-2}=\frac{1}{3^2}=\frac{1}{9}\] that is how you are "supposed to do it"

OpenStudy (anonymous):

now get out your fancy ti 83 and type in \[\log(1\div 9)\div \log(3)\] and out will pop -2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so if you wish to cheat your teacher and find \[\log_4(\frac{1}{8})\] without saying \[4^{\frac{3}{2}}=8\] then just plug in \[\log(1\div8)\div \log(4)\] and out will pop \[\frac{3}{2}\] try it!

OpenStudy (anonymous):

always happy to help some one do it the easy way.